# Dynamics and modulation of cosmic ray modified magnetosonic waves in a galactic gaseous rotating plasma

### v2

@@ I INTRODUCTION
In our Galaxy, interstellar medium (ISM) clouds are highly inhomogeneous and comprise various phases of gas clouds, such as a hot phase, a warm phase, diffusive clouds, molecular clouds, etc. These components of the ISM cloud exhibit a similar energy density distributionkuwabara2020dynamics . It is well-known that the galactic cosmic rays significantly contribute to the evolution and structure formation of ISM clouds. Galactic cosmic rays parker1969galactic ; padovani2020impact are highly energetic particles ejected from astrophysical bodies outside our solar system, such as supernova remnants, planetary disks, the Sun, etc., which numerously interact with different components of the ISM. Cosmic rays play an important role in the charge of dust grains kalvans2016temperature , gas ionization yusef2007cosmic , and energy transfer gaches2019astrochemical in the interstellar medium. They interact with cosmic dust grains via different physical mechanisms, leading to the generation of different wave modes and instabilities in the media. They are major contributors to the ionization of molecular clouds (MCs) core padovani2011effects and the ionization of molecule hydrogen ($H_{2}$) in diffuse interstellar clouds indriolo2009implications . They transfer energy to the dust grains through heating the dust grains in cosmic ray-plasma interactions.
@@
The nonlinearities in the plasma contribute to energy localization, leading to the evaluation of various kinds of nonlinear coherent structures, like solitons, shocks, etc., which have immense application in space as well as astrophysical plasma environmentsgoertz1989dusty ; havnes1992charged ; horanyi1986effects . The evolutions of these nonlinear coherent structures are characterized by various well-known nonlinear partial differential equations, such as Korteweg-de Vries (KdV) and KdV-like equations wazwaz2017abundant , the Zakharov-Kuznetsov (ZK) and ZK-like equations, and Kadomtsev-Petviashvili (KP) and KP-like equations chen2018special ; el2016generation . Another crucial equation that controls the motion of nonlinear structures is the nonlinear Schrödinger (NLS) equationmihalache2017multidimensional . A rogue waveakhmediev2009waves ; muller2005rogue is one of the viable rational solutions of the NLS equationsharma2024modulation ; chen2015vector . Such waves are often referred to as monster, violent, extreme, or giant waves that have a few times higher amplitude than solitary waves and are observed in coastal waters kharif2008rogue , fibre optic and optical systems solli2007optical , Bose-Einstein condensates bludov2009matter , atmosphere stenflo2010rogue and astrophysical environments sabry2012freak . There has been increasing interest in investigating the behaviour of magnetosonic waves and the associated coherent structures like solitons, shocks, rogue waves, etc., in different plasma environmentsrahim2019magnetosonic ; hager2023magnetosonic ; el2012rogue ; moslem2011surface ; akbari2014electrostatic .
@@
In recent times, researchers have devoted significant attention to the role of cosmic rays in various wave modes in linear regimes and their stability/instabilities. In this direction, Turi and Misra turi2022magnetohydrodynamic have addressed the impact of cosmic rays on the propagation of MHD waves as well as their instability (Jeans) for a self-gravitating galactic gaseous cloud. In another work, Boro and Prajapati boro2023cosmic have studied the Jeans instability for the molecular dusty plasma environment of molecular clouds, including cosmic rays. Further, various previous studies suggest that the Coriolis force plays an important role in various plasma phenomena in different plasma environments, like rotating plasma in the laboratory as well as astrophysical plasmaturi2022magnetohydrodynamic ; hager2023magnetosonic . To the best of our knowledge, the influence of cosmic rays on the dynamics of nonlinear coherent structures in ISM clouds, under the effect of Coriolis force, has not yet been addressed. Thus, in the current work, we aim to explore the impact of cosmic rays on linear and nonlinear magnetosonic shocks, solitons, and rogue waves in a dissipative rotating magnetoplasma consisting of thermal gas and cosmic rays. The present model is suitable and well justified in real physical situations like ISM clouds in spiral galaxies. For a fairly good approximation in studying the structures and evolution of this system, a hydrodynamic description including the cosmic rays has been considered.
@@
This present work is organised as follows: in Sec. II, we describe the basic equations for the MHD waves. In Sec. III, linear properties of the magnetosonic waves are discussed through the linear dispersion relation. Sec. IV includes nonlinear analysis, where we derive a Korteweg–de Vries–Burgers (KdVB) equation for the magnetosonic waves and analyse the analytical and numerical behaviour of the equation. Sec. V is devoted to modulational instability and rogue wave solutions for carrier wave frequencies much smaller than the wave frequency. Finally, we summarize our main findings in Sec. VI.
Figure 1: The real part $\omega_{r}\equiv\omega_{r}(k)$ and the damping rate $\gamma\equiv\gamma_{(}k)$ of the dispersion relation (15) are displayed for different values of the parameters as mentioned in the legends. Here, ($\theta$, $\kappa$, $\eta$)= ($\pi/3,0.3,0.03)$ in the subplots (a) & (b) and $(\theta,\Omega_{0},C_{g},C_{c})=(\pi/3,0.9,0.2,0.4)$ in the subplot (c). The other parameter values are $\rho_{0}\sim 10^{-21}Kg/m^{3}$, $B_{0}\sim 1$ nT, $P_{g0},P_{c0}\sim 10^{-13}N/m^{2}$, $\eta\sim 0.4(V_{A}^{2}/\omega_{ci})s^{-1}$ and $\kappa\sim 0.06(V_{A}^{2}/\omega_{ci})s^{-1}$.
@@ II BASIC EQUATIONS OF THE PROBLEM
The typical plasma environment of the interstellar medium, composed of ionized thermal gas and cosmic rays, has been considered for the study of the propagation of magnetosonic waves. Here, the thermal ionized gas is considered polytropic, having no diffusion in any direction, and cosmic rays are assumed to be gas with negligible density but contributing significantly to the pressure of the system bonanno2008generation ; turi2022magnetohydrodynamic . The plasma is supposed to be immersed in a uniform external magnetic field $\mathbf{B}=B(x,t)\hat{e}_{z}$, where $B(x,t)$ is the static external magnetic field strength and $\hat{e}_{z}$ denotes the unit vector along the $z$ direction. The environment of the interstellar medium is considered to be homogeneous, fully ionized, and highly conductive fluid with a higher Reynold’s number. The plasma system is assumed to rotate slowly with a rotational frequency $\Omega$ around an axis lying in the $xz$-plane. The basic evolution equations describing the dynamics of magnetosonic waves are given as turi2022magnetohydrodynamic
@@
The above set of Eqs. (6)-(12) is made normalized by normalising the variables and parameters as follows: $\rho\rightarrow\rho/\rho_{0}$, $B\rightarrow B/B_{0}$, $t\rightarrow t\omega_{ci}$, $v_{x}\rightarrow v_{x}/V_{A}$, $v_{y}\rightarrow v_{y}/V_{A}$, $v_{z}\rightarrow v_{z}/V_{A}$, $x\rightarrow x\omega_{ci}/V_{A}$, $\Omega\rightarrow\Omega/\omega_{ci}$, $P_{g}\rightarrow P_{g}/\rho_{0}V_{A}^{2}$, $P_{c}\rightarrow P_{c}/\rho_{0}V_{A}^{2}$, $\eta\rightarrow\eta\omega_{ci}/V_{A}^{2}$, $\kappa\rightarrow\kappa\frac{\omega_{ci}}{V_{A}^{2}}$, where, $\rho_{0}$ is the equilibrium density, $B_{0}$ is the equilibrium magnetic field strength, and $\omega_{ci}=\frac{eB_{0}}{m_{i}}$ is the ion cyclotron frequency, $V_{A}=\frac{B_{0}}{\sqrt{\mu_{0}\rho_{0}}}$ is the Alfvén velocity. Furthermore, our analysis is suitable based on real magnetoplasma systems, which are desired to be relevant in spiral galaxies gliddon1966gravitational ; turi2022magnetohydrodynamic . Consequently, for numerical demonstration, we choose the typical consistent parameter values as $B_{0}\sim 1$ nT, $\rho_{0}\sim 10^{-21}$ Kg/m3, $P_{g0}\sim 10^{-13}$ N/m2, $P_{c0}\sim 10^{-13}$/m2 and the other parameters are taken as $\eta\sim 10^{-1}V_{A}^{2}/\omega_{ci}$, $\kappa\sim 10^{-2}V_{A}^{2}/\omega_{ci}$. In this regime, the ion gyro-frequency is typically scaled as $\omega_{ci}\sim 10^{-1}$ s-1. The numerical value of the rotational frequency is taken to be much lower than the ion gyro-frequency $\omega_{ci}$ abdikian2022drift .
Figure 2: One-soliton profiles $B~{}\text{vs}~{}\zeta$ are plotted for different values of the parameters as mentioned in the legends. Here, $(C_{g},C_{c})$= $(0.2,0.4)$ and $(\Omega_{0},\theta)=(0.9,\pi/3)$ in the subplots (a) and (b), respectively, and $v_{0}=0.1$. The other parameters remain unchanged as in Fig. 1.
@@ III LINEAR ANALYSIS
We study the propagation of magnetosonic plane waves in this plasma system along the $x$-direction. To proceed, we linearize the set of Eqs. (6)-(12) about the equilibrium state by assuming the plasma is uniform with constant density and zero velocity. Dividing the physical quantity $f$ into the equilibrium ($f_{0}$) and perturbed ($f_{1}$) parts as $f=f_{0}+f_{1}$, where $f=(\rho,B,v_{x},v_{y},v_{z},P_{g},P_{c})$, $f_{0}=(1,1,0,0,0,P_{g0},P_{c0})$ and $f_{1}=(\rho_{1},B_{1},v_{x1},v_{y1},v_{z1},P_{g1},P_{c1})$, we obtain a linearized set of equations from Eqs. (6)-(12).
@@
where $C_{g}(=\sqrt{\gamma_{g}P_{g0}})$ and $C_{c}(=\sqrt{\gamma_{c}P_{c0}})$ are the normalized form (normalized by $V_{A}$) of the speed of thermal gas and cosmic rays, respectively. It is worth mentioning that the denormalization of the above quantities yields $C_{g}=\sqrt{\frac{\gamma_{g}P_{g0}}{\rho_{0}}}$ and $C_{c}=\sqrt{\frac{\gamma_{c}P_{c0}}{\rho_{0}}}$. The dispersion relation (III) agrees with a similar result as in Ref. turi2022magnetohydrodynamic in the absence of magnetic resistivity ($\eta$) and cosmic ray diffusivity ($\kappa$). Due to the presence of the parameters $\eta$ and $\kappa$ in the dispersion relation, $\omega$ and $k$ become complex. The dispersion is also influenced by the rotational frequency ($\Omega_{0}$) and the angle $\theta$ related to the rotation of the fluid in the $xz$-plane. If the axis of rotation is taken along $x$-axis (i.e. $\theta=\pi/2$), the dispersion relation (III) modified to
@@
where the effects of the Coriolis force disappear due to the rotation of the plasma along the $x$-axis, however, the influences of magnetic resistivity and cosmic ray diffusivity have been retained. The wave becomes dispersionless in long-wavelength limits (i.e. when $k\ll 1$) and/or when the effects of magnetic resistivity and cosmic ray diffusivity weaken and the dispersion reduces to $\omega=ak$ where $a=(1+C_{g}^{2}+C_{c}^{2})^{1/2}$. In the case of when the axis of rotation is taken along the $z$-axis (i.e. $\theta=0$), the dispersion relation (III) reduces to
@@
The dispersion (15) includes the effect of the Coriolis force due to the rotation of the plasma. The wave becomes unstable because of the presence of cosmic rays and magnetic diffusion, i.e., it can be either damped or anti-damped. To obtain the growth rate (or damping) we separate the real and imaginary parts of the wave frequency as $\omega=\omega_{r}+i\gamma$ with the assumption $|\gamma|$, $|\kappa|$, $|\eta|\ll\omega_{r}$, $\frac{\omega}{\omega+i\kappa k^{2}}\approx(1-i\frac{\kappa k^{2}}{\omega_{1}})$ and $\frac{\omega}{\omega+i\eta k^{2}}\approx(1-i\frac{\eta k^{2}}{\omega_{1}})$, where $\omega_{1}$ is the solution of Eq. (15) at $\eta=\kappa=0$. We found that the magnetosonic wave gets damped, and the absolute value of damping is estimated to be $|\gamma|\sim\frac{1}{2}(C_{c}^{2}\kappa+\eta)\frac{k^{4}}{\omega_{1}^{2}}$ and the real part is $\omega_{r}=\omega_{1}$.
@@
To study the dispersion and damping characteristic of the magnetosonic wave, we take numerical results from Eq. (15) displayed in Fig. 1. The plots of the real part $\omega_{r}=\omega_{r}(k)$ are shown in the subplot (a) for different values of the parameters $\Omega_{0}$, $C_{g}$ and $C_{c}$ associated with the rotational frequency, thermal pressure and cosmic ray pressure, respectively. It is found that initially, for $k\ll 1$, the profile of $\omega_{r}$ remains parallel to the axis of wave number, i.e., the phase velocity approaches a constant value, as we discussed earlier. Beyond the limit $k\ll 1$, the frequency of the wave increases with increasing values of $k$. The profiles ($\omega_{r}$) are plotted for four sets of values of $\Omega_{0}$, $C_{g}$ and $C_{c}$. It has been observed from the profiles that an increase in either of these three parameters, keeping the values of alternate parameters fixed, leads to an enhancement of the wave frequency. However, in the limit $k\ll 1$, the increases in wave frequencies are found to be negligible for increasing values of $C_{g}$ and $C_{c}$, beyond the limit, the increments are transpicuous. In the subplots (b) and (c) of Fig. 1, the profiles of estimated damping rate $\gamma=\gamma(k)$ are plotted for different values of $\Omega_{0}$, $C_{g}$, $C_{c}$ and $\kappa$, $\eta$, respectively. The figures show that the absolute damping rate is negligible in the limit $k\ll 1$ but the damping grows with increasing values of $k$. Subplot (b) depicts that an increase in the values of $\Omega_{0}$ and $C_{g}$ decreases the absolute damping rate, whereas the wave damping increases with increasing value of the parameter $C_{c}$. On the other hand, from the subplot (c), the significant growth in the damping rate has been noticed with the increasing values of $\eta$ and $\kappa$, which leads to unstable behaviour in the system. Magnetic resistivity $\eta$ and cosmic ray diffusivity $\kappa$ do not contribute to the real part of the wave frequency but influence the wave damping. However, subplots (a) and (b) show that the absolute damping $\gamma$ reduces as the rotating frequency increases, but the real part of the wave frequency increases. So, it is interesting to note that the damping due to magnetic resistivity and cosmic ray diffusivity becomes insignificant with an increase in rotating frequency.
@@
Figure 3: Two-soliton profiles $B~{}\text{vs}~{}\xi$ are displayed for different values of $\tau$. Here, $(\Omega_{0},\theta,C_{g},C_{c})=(0.7,\pi/4,0.2,0.4)$, $k_{1}=0.5$ and $k_{2}=1.2$. The other parameters remain unchanged as in Fig. 1.

### v3

@@ I INTRODUCTION
In our Galaxy, interstellar medium (ISM) clouds are highly inhomogeneous and comprise various phases of gas clouds, such as a hot phase, a warm phase, diffusive clouds and molecular clouds. These components of the ISM cloud exhibit a similar energy density distributionkuwabara2020dynamics . It is well-known that the galactic cosmic rays significantly contribute to the evolution and structure formation of ISM clouds. Galactic cosmic rays parker1969galactic ; padovani2020impact are highly energetic particles ejected from astrophysical bodies outside our solar system, such as supernova remnants, planetary disks and the Sun, which numerously interact with different components of the ISM. Cosmic rays play an important role in the charge of dust grains kalvans2016temperature , gas ionization yusef2007cosmic , and energy transfer gaches2019astrochemical in the interstellar medium. They interact with cosmic dust grains via different physical mechanisms, leading to the generation of different wave modes and instabilities in the media. They are major contributors to the ionization of molecular clouds (MCs) core padovani2011effects and the ionization of molecule hydrogen ($H_{2}$) in diffuse interstellar clouds indriolo2009implications . They transfer energy to the dust grains through heating the dust grains in cosmic ray-plasma interactions.
@@
The nonlinearities in the plasma contribute to energy localization, leading to the evaluation of various types of nonlinear coherent structures, viz., solitons and shocks, which have immense application in space as well as astrophysical plasma environmentsgoertz1989dusty ; havnes1992charged ; horanyi1986effects . The evolutions of these nonlinear coherent structures are characterized by various well-known nonlinear partial differential equations, such as Korteweg-de Vries (KdV), Korteweg–de Vries–Burgers (KdVB), Zakharov-Kuznetsov (ZK), Kadomtsev-Petviashvili (KP) equations and their familieswazwaz2017abundant ; chen2018special ; el2016generation . Another crucial equation that controls the motion of nonlinear structures is the nonlinear Schrödinger (NLS) equationmihalache2017multidimensional . A rogue waveakhmediev2009waves ; muller2005rogue is one of the viable rational solutions of the NLS equationsharma2024modulation ; chen2015vector . Such waves are often referred to as monster, violent, extreme, or giant waves that have a few times higher amplitude than solitary waves and are observed in coastal waters kharif2008rogue , fibre optic and optical medium solli2007optical , Bose-Einstein condensates bludov2009matter , atmosphere stenflo2010rogue and astrophysical environments sabry2012freak . There has been increasing interest in investigating the behaviour of magnetosonic waves and the associated coherent structures, viz., solitons, shocks and rogue waves in different plasma environmentsrahim2019magnetosonic ; hager2023magnetosonic ; el2012rogue ; moslem2011surface ; akbari2014electrostatic .
@@
In recent times, researchers have devoted significant attention to the role of cosmic rays in various wave modes in linear regimes and their stability/instabilities. To mention a few, in a recent work, Marcowith et al. marcowith2021cosmic have addressed Cosmic ray-driven streaming instability in different space and astrophysical plasmas. Zhou et al. zhou2024effects have studied the impact of cosmic rays on the dynamics of solar wind in the outer heliosphere, considering three components of magnetohydrodynamic (MHD) plasma consisting of the solar wind, interstellar neutral atoms, and cosmic rays. The study shows that cosmic rays may decrease the speed of the solar wind shocks and affect the shock structure. In this direction, Turi and Misra turi2022magnetohydrodynamic have addressed the impact of cosmic rays on the propagation of MHD waves as well as their instability (Jeans) for a self-gravitating galactic gaseous cloud. Mansuri et al. mansuri2023evolution have theoretically studied the Jeans instability considering cosmic rays in a radiative quantum plasma, including Hall current and Coriolis force. It is reported that cosmic rays significantly alter the excitation of Jeans instability and affect the instability region as well as the instability growth rate. In another work, Boro et al. boro2023cosmic have studied the Jeans instability for the molecular dusty plasma environment of molecular clouds, including cosmic rays. Recently, some new exciting features of low-frequency MHD waves and associated instability (Jeans) consisting of superthermal gas and cosmic rays in an astrophysical plasma environment have been addressed by Boro and Prajapati boro2024suprathermal . Further, various previous studies suggest that the Coriolis force plays an important role in various plasma phenomena in different plasma environments, such as rotating plasma in the laboratory as well as astrophysical plasmaturi2022magnetohydrodynamic ; hager2023magnetosonic .
Existing literature reveals that there is still some lack of studies in the direction of theoretical investigation of magnetosonic waves in the ISM of spiral galaxies, exploring the effect of cosmic rays’ interaction with ionized gas in such dissipative rotating magnetoplasma. The main objective of the present study is to advance the theory of magnetosonic wave dynamics in the ISM of spiral galaxies, consisting of thermally ionized gas and cosmic rays. The cosmic rays are assumed to interact with the thermally ionized gas in a fluid-fluid approach and modify the existing pressure law of the plasma. We aim to explore the impact of cosmic rays, in terms of the parameters related to the cosmic ray pressure and diffusion, on both the linear and nonlinear magnetosonic waves. Toward this objective, we use the basic MHD equations including the cosmic ray modified pressure law for a fairly good approximation in studying the structures and evolution of the magnetosonic waves. The model is suitable and well justified in real physical situations of ISM clouds in spiral galaxies.
@@
The manuscript is organised as follows: in Sec. II, we describe the basic equations for the MHD waves. In Sec. III, linear properties of the magnetosonic waves are discussed through the linear dispersion relation. Sec. IV includes nonlinear analysis, where we derive a Korteweg–de Vries–Burgers (KdVB) equation for the magnetosonic waves and analyse the analytical and numerical behaviour of the equation. Sec. V is devoted to modulational instability and rogue wave solutions for carrier wave frequencies much smaller than the wave frequency. A sensitivity analysis for the plasma parameters is performed in Sec. VI. Finally, we summarize our main findings in Sec. VII.
Figure 1: The real part $\omega_{r}\equiv\omega_{r}(k)$ and the damping rate $\gamma\equiv\gamma(k)$ of the dispersion relation (15) are displayed for different values of the parameters as mentioned in the legends. The profiles show that, for $k\ll 1$, the wave frequency ($\omega_{r}$) remains constant and the wave damping ($\gamma$) is negligible. Beyond the limit, both increase for increasing values of $k$. The damping occurs due to the effects of the parameters $C_{g}$ and $C_{c}$ related to thermal and cosmic ray pressure, rotational frequency ($\Omega_{0}$), cosmic ray diffusivity ($\kappa$) and magnetic resistivity ($\eta$). Here, ($\theta$, $\kappa$, $\eta$)= ($\pi/3,0.3,0.03)$ in the subplots (a) & (b) and $(\theta,\Omega_{0},C_{g},C_{c})=(\pi/3,0.9,0.2,0.4)$ in the subplot (c). These dimensionless typical plasma parameters for the ISM of spiral galaxies are chosen from Table I.
@@ II BASIC EQUATIONS OF THE PROBLEM
The typical plasma environment of the interstellar medium, composed of ionized thermal gas and cosmic rays, has been considered to study the propagation of magnetosonic waves. Here, the thermal ionized gas is considered polytropic, having no diffusion in any direction, and cosmic rays are assumed to be gas with negligible density but contributing significantly to modify the pressure of the plasma environment bonanno2008generation ; turi2022magnetohydrodynamic . The plasma is supposed to be immersed in a uniform external magnetic field $\mathbf{B}=B(x,t)\hat{e}_{z}$, where $B(x,t)$ is the static external magnetic field strength and $\hat{e}_{z}$ denotes the unit vector along the $z$ direction. The interstellar medium’s environment is considered to be homogeneous, fully ionized and conducting fluid. The plasma is assumed to rotate slowly with a rotational frequency $\Omega$ around an axis lying in the $xz$-plane. The basic evolution equations describing the dynamics of magnetosonic waves are given as turi2022magnetohydrodynamic ; Turi_2024
@@
The above set of Eqs. (6)-(12) is made normalized by normalising the variables and parameters as follows: $\rho\rightarrow\rho/\rho_{0}$, $B\rightarrow B/B_{0}$, $t\rightarrow t\omega_{ci}$, $v_{x}\rightarrow v_{x}/V_{A}$, $v_{y}\rightarrow v_{y}/V_{A}$, $v_{z}\rightarrow v_{z}/V_{A}$, $x\rightarrow x\omega_{ci}/V_{A}$, $\Omega\rightarrow\Omega/\omega_{ci}$, $P_{g}\rightarrow P_{g}/\rho_{0}V_{A}^{2}$, $P_{c}\rightarrow P_{c}/\rho_{0}V_{A}^{2}$, $\eta\rightarrow\eta\omega_{ci}/V_{A}^{2}$, $\kappa\rightarrow\kappa\omega_{ci}/V_{A}^{2}$, where, $\rho_{0}$ is the equilibrium density, $B_{0}$ is the equilibrium magnetic field strength, and $\omega_{ci}=\frac{eB_{0}}{m_{i}}$ is the ion cyclotron frequency, $V_{A}=\frac{B_{0}}{\sqrt{\mu_{0}\rho_{0}}}$ is the Alfvén velocity. Furthermore, our analysis is suitable based on real magnetoplasma environments, which are desired to be relevant in spiral galaxies gliddon1966gravitational ; turi2022magnetohydrodynamic . Consequently, for numerical demonstration, we choose the typical consistent parameter values as given in Table I. In this regime, the ion gyro-frequency is typically scaled as $\omega_{ci}\sim 10^{-1}$ s-1. The plasma is assumed to be rotated with small rotational frequency, i.e. $\Omega_{0}<1$ allowing us to ignore the second- and higher-order terms of $\Omega_{0}$. Hence, we neglect the effect of centrifugal force $\Omega_{0}\times(\Omega_{0}\times r)$abdikian2022drift ; abdikian2021supernonlinear . Denormalization of $\Omega_{0}$ implies $\Omega_{0}/\omega_{ci}\ll 1$. Thus, the numerical value of the rotational frequency is taken to be much lower than the ion gyro-frequency $\omega_{ci}$.
Physical quantity Symbol Value Units
Fluid density $\rho_{0}$ $2\times 10^{-21}$ Kg/m3
Magnetic field $B_{0}$ $1\times 10^{-9}$ T
Thermal pressure $P_{g0}$ $2.95\times 10^{-13}$ N/m2
Cosmic ray pressure $P_{c0}$ $2.14\times 10^{-13}$ N/m2
Electron charge $e$ $1.6\times 10^{-19}$ C
Ion mass $m_{i}$ $1.67\times 10^{-27}$ kg
Thermal speed $C_{g}$ $1.39\times 10^{4}$ m s-1
Cosmic ray speed $C_{c}$ $1.19\times 10^{4}$ m s-1
Rotational frequency $\Omega_{0}$ $0.076$ s-1
Adiabetic index of thermal gas $\gamma_{g}$ 4/3 Unit less
Adiabetic index of cosmic rays $\gamma_{c}$ 4/3 Unit less
Permeability of free space $\mu_{0}$ $1.26\times 10^{-6}$ H/m
Magnetic resistivity $\eta$ $8.28\times 10^{7}$ m2 s-1
Cosmic ray diffusivity $\kappa$ $3.72\times 10^{9}$ m2 s-1
Table 1: Values of the typical plasma parameters of spiral galaxies.gliddon1966gravitational ; turi2022magnetohydrodynamic
Figure 2: The typical one-soliton profiles $B~{}\text{vs}~{}\zeta$ are plotted [see Eq. (26)] for different values of the parameters as mentioned in the legends. The wave amplitude remains unchanged as the rotating frequency $\Omega_{0}$ increases, however, it decreases due to the increasing value of the angle of rotation ($\theta$) and parameters related to thermal pressure ($C_{g}$) and cosmic ray pressure ($C_{c}$). The width of the waves decreases due to the increasing effect of $\Omega_{0}$ and $\theta$, but it becomes wider as $\theta$, $C_{g}$ and $C_{c}$ increase. Here, $(C_{g},C_{c})$= $(0.2,0.4)$ and $(\Omega_{0},\theta)=(0.9,\pi/3)$ in the subplots (a) and (b), respectively, and $v_{0}=0.1$. The other parameters remain unchanged as in Fig. 1.
@@ III LINEAR ANALYSIS
We study the propagation of magnetosonic plane waves in this plasma medium along the $x$-direction. To proceed, we linearize the set of Eqs. (6)-(12) about the equilibrium state, assuming the plasma is uniform with constant density and zero velocity. Dividing the physical quantity $f$ into the equilibrium ($f_{0}$) and perturbed ($f_{1}$) parts as $f=f_{0}+f_{1}$, where $f=(\rho,B,v_{x},v_{y},v_{z},P_{g},P_{c})$, $f_{0}=(1,1,0,0,0,P_{g0},P_{c0})$ and $f_{1}=(\rho_{1},B_{1},v_{x1},v_{y1},v_{z1},P_{g1},P_{c1})$, we obtain a linearized set of equations from Eqs. (6)-(12).
@@
where $C_{g}(=\sqrt{\gamma_{g}P_{g0}})$ and $C_{c}(=\sqrt{\gamma_{c}P_{c0}})$ are the normalized form (normalized by $V_{A}$) of the speed of thermal gas and cosmic rays, respectively. It is worth mentioning that the denormalization of the above quantities yields $C_{g}=\sqrt{\frac{\gamma_{g}P_{g0}}{\rho_{0}}}$ and $C_{c}=\sqrt{\frac{\gamma_{c}P_{c0}}{\rho_{0}}}$. Due to the presence of the parameters $\eta$ and $\kappa$ in the dispersion relation, $\omega$ and $k$ become complex. The dispersion is also influenced by the rotational frequency ($\Omega_{0}$) and the angle $\theta$ related to the rotation of the fluid in the $xz$-plane. If the axis of rotation is taken along $x$-axis (i.e. $\theta=\pi/2$), the dispersion relation (III) modified to
@@
where the effects of the Coriolis force disappear due to the rotation of the plasma along the $x$-axis, however, the influences of magnetic resistivity and cosmic ray diffusivity have been retained. The wave becomes dispersionless in long-wavelength limits (i.e. when $k\ll 1$) and/or when the effects of magnetic resistivity and cosmic ray diffusivity are weaken. Consequently, the dispersion reduces to $\omega=ak$ where $a=(1+C_{g}^{2}+C_{c}^{2})^{1/2}$. In the case of when the axis of rotation is taken along the $z$-axis (i.e. $\theta=0$), the dispersion relation (III) reduces to
@@
The dispersion (15) includes the effect of the Coriolis force due to the rotation of the plasma. Assuming the plasma to be a highly conducting fluid, if the effect of magnetic resistivity is ignored, the above-normalized dispersion relation (15) reduces to the dispersion relation as obtained by Turi and Misraturi2022magnetohydrodynamic upon denormalization, in the absence of self-gravitation. Further, disregarding the effect of Coriolis force due to the rotation of the fluid, cosmic ray pressure, and cosmic ray diffusion, one can recover the typical magnetosonic mode with phase velocity $\omega/k=(C_{g}^{2}+V_{A}^{2})^{1/2}$ as given in Bittencourt bittencourt2013fundamentals . Due to the effect of cosmic ray diffusivity and magnetic resistivity, the wave becomes unstable, i.e., it can be either damped or anti-damped. To obtain the growth rate (or damping), we separate the real and imaginary parts of the wave frequency as $\omega=\omega_{r}+i\gamma$, assuming $|\gamma|$, $|\kappa|$, $|\eta|\ll\omega_{r}$, $\frac{\omega}{\omega+i\kappa k^{2}}\approx(1-i\frac{\kappa k^{2}}{\omega_{1}})$ and $\frac{\omega}{\omega+i\eta k^{2}}\approx(1-i\frac{\eta k^{2}}{\omega_{1}})$, where $\omega_{1}$ is the solution of Eq. (15) at $\eta=\kappa=0$. We found that the magnetosonic wave gets damped, and the absolute value of damping is estimated to be $|\gamma|\sim\frac{1}{2}(C_{c}^{2}\kappa+\eta)\frac{k^{4}}{\omega_{1}^{2}}$ and the real part is $\omega_{r}=\omega_{1}$. The dispersion relation (15) includes complex parts due to the contribution of $\eta$ and $\kappa$ to the wave mode. In the context of spiral galaxies, a significant contribution of $\eta$ and $\kappa$ causes damping into the waves, resulting in wave energy decay, hence transferring energy to the ISM. This transformation of wave energy to thermal energy contributes to heating the ISM of spiral galaxies.
@@
To study the dispersion and damping characteristic of the magnetosonic wave, we take numerical results from Eq. (15) displayed in Fig. 1. The plots of the real part $\omega_{r}=\omega_{r}(k)$ are shown in the subplot (a) for different values of the parameters $\Omega_{0}$, $C_{g}$ and $C_{c}$ associated with the rotational frequency, thermal pressure and cosmic ray pressure, respectively. It is found that initially, for $k\ll 1$, the profile of $\omega_{r}$ remains parallel to the axis of wave number, i.e., the phase velocity approaches a constant value, as we discussed earlier. Beyond the limit $k\ll 1$, the frequency of the wave increases with increasing values of $k$. The profiles ($\omega_{r}$) are plotted for four sets of values of $\Omega_{0}$, $C_{g}$ and $C_{c}$. It has been observed from the profiles that an increase in either of these three parameters, keeping the values of alternate parameters fixed, leads to an enhancement of the wave frequency. However, in the limit $k\ll 1$, the increases in wave frequencies are found to be negligible for increasing values of $C_{g}$ and $C_{c}$, beyond the limit, the increments are transpicuous. In the subplots (b) and (c) of Fig. 1, the profiles of estimated damping rate $\gamma=\gamma(k)$ are plotted for different values of $\Omega_{0}$, $C_{g}$, $C_{c}$ and $\kappa$, $\eta$, respectively. The figures show that the absolute damping rate is negligible in the limit $k\ll 1$, but the damping grows with increasing values of $k$. Subplot (b) depicts that an increase in the values of $\Omega_{0}$ and $C_{g}$ decreases the absolute damping rate, whereas the wave damping increases with increasing value of the parameter $C_{c}$. On the other hand, from the subplot (c), the significant growth in the damping rate has been noticed with the increasing values of $\eta$ and $\kappa$, which leads to unstable behaviour on wave mode. Magnetic resistivity $\eta$ and cosmic ray diffusivity $\kappa$ do not contribute to the real part of the wave frequency but influence the wave damping. However, subplots (a) and (b) show that the absolute damping $\gamma$ reduces as the rotating frequency increases, but the real part of the wave frequency increases. So, it is interesting to observe that the damping due to magnetic resistivity and cosmic ray diffusivity becomes insignificant with an increase in rotating frequency.
@@
Figure 3: Two-soliton profiles $B~{}\text{vs}~{}\xi$ are displayed [see Eq. (29)] for different values of $\tau$ showing the two solitons with different amplitudes collide and exchange energy and phase. For $\tau<0$ [see subplots (a) and (b)], the soliton with higher amplitude stays behind the lesser one. At $\tau=0$ [see subplot (c)] two solitons collide and become one-soliton, after that, the soliton with higher amplitude moves faster than the lesser one [see subplots (d) and (e)]. Here, $(\Omega_{0},\theta,C_{g},C_{c})=(0.7,\pi/4,0.2,0.4)$, $k_{1}=0.5$ and $k_{2}=1.2$. The other parameters remain unchanged as in Fig. 1.